OpenStudy (anonymous):

4(2x - 1) - 3 = 9(x - 4)

OpenStudy (acxbox22):

distribute the 4 and the 9 first then solve for x

OpenStudy (anonymous):

4(2x-1)-3=9(x-4) 8x-4-3=9x-4 x=-3